# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v1

@@ VIII Summary
In this comprehensive review, we have detailed the advancements and capabilities of the ABACUS software, a powerful open-source platform for first-principles electronic structure calculations and molecular dynamics simulations. ABACUS stands out for its versatility, supporting both plane-wave and numerical atomic orbital basis sets, and its compatibility with a range of electronic structure methods.
@@
The methods based on the plane-wave basis in ABACUS have been continuously developed. Stochastic DFT has been implemented, offering an alternative approach to handle large systems with improved computational efficiency. Orbital-Free DFT has also been included, with different kinetic energy density functionals and optimization methods available. For calculations using the numerical atomic orbital basis, ABACUS provides efficient implementations. The methods such as hybrid functionals, DFT+U, and Real-Time TDDFT have been extended to this basis set. The integration of machine learning techniques in the form of the DeePKS method has shown great potential in achieving high accuracy at a reduced computational cost. These methods have been benchmarked and applied to various systems, demonstrating their accuracy and effectiveness.
@@
ABACUS has actively participated in the OpenLAM project, generating high-precision first-principles data for different scenarios, including semiconductors, alloys, ferroelectric, and superhydrides. The APNS project has been initiated to ensure the reliability and accuracy of ABACUS. Systematic tests for pseudopotentials and numerical atomic orbitals have been conducted, providing a solid foundation for the software’s performance. The software’s ability to interface with other packages, such as DeePKS-kit, DeePMD, DP-GEN, DeepH, DeePTB, PyATB, Hefei-NAMD, ASE, Phonopy, TB2J, and Wannier90, further expands its functionality and applicability in multi-scale simulations and analysis.
@@
In conclusion, ABACUS has established itself as a robust and flexible platform for electronic structure calculations, poised to meet the challenges of materials science, chemistry, and physics in the AI era. Its ongoing development and integration with AI technologies promise to further enhance its capabilities, making it an indispensable tool for researchers and scientists in the field.
@@ Acknowledgements.
We thank Han Wang and Weinan E for many helpful discussions. M. C. gratefully acknowledges AI for Science Institute, Beijing (AISI), as well as funding supports from the National Natural Science Foundation of China (grant no. 12122401,12074007,12135002). S. X. gratefully acknowledges funding support from the National Natural Science Foundation of China (grant no. 52273223), DP Technology Corporation (grant no. 2021110016001141), School of Materials Science and Engineering at Peking University, and the AI for Science Institute, Beijing (AISI). S. L. acknowledges the supports from National Natural Science Foundation of China (92370104) and Natural Science Foundation of Zhejiang Province (2022XHSJJ006). The computational resource is provided by Westlake HPC Center. M. L. gratefully acknowledges the financial support of the National Natural Science Foundation of China (No. 52276212), National Key Research and Development Program of China (No. 2022YFB3803600). Q. Ou acknowledges research grants from China Petroleum &\&& Chemical Corp (funding number 124014). T.W. acknowledges support by The University of Hong Kong (HKU) via seed funds (2201100392, 2409100597). Y.X. acknowledges support by the National Key Basic Research and Development Program of China (grant nos. 2024YFA1409100 and 2023YFA1406400), the National Natural Science Foundation of China (grants nos. 12334003, 12421004, and 12361141826), and the National Science Fund for Distinguished Young Scholars (grant no. 12025405).

### v3

@@ VII Summary
In this review, we have detailed the recent advances in the ABACUS package, which is an open-source platform for first-principles electronic structure calculations and molecular dynamics simulations. ABACUS supports both PW and NAO basis sets, and is compatible with a range of electronic structure methods.
@@
On one hand, built on a plane-wave (PW) basis, ABACUS supports iterative methods with parallel algorithms for diagonalizing the Kohn-Sham equation. Additionally, the stochastic DFT method has been implemented, offering an alternative approach to address extremely high-temperature systems with substantially improved computational efficiency. Furthermore, OFDFT has been incorporated, featuring diverse kinetic energy density functionals, including machine learning-based ones, and optimization methods tailored for large-scale materials systems. On the other hand, by using the NAO basis set, ABACUS provides efficient solutions to the Kohn-Sham equation, featuring methods such as hybrid functional, DeePKS, and DFT+U, etc. In addition, real-time TDDFT has been implemented. These methods have been benchmarked and applied to various systems, demonstrating their accuracy and effectiveness.
@@
ABACUS has been used to generate a large amount of first-principles data for training machine-learning-based potentials. The APNS project has been initiated to ensure the reliability and accuracy of ABACUS, where systematic tests for pseudopotentials and numerical atomic orbitals have been conducted, providing a solid foundation for the software’s performance. ABACUS interacts with several packages, such as DeePKS-kit, DeePMD-kit, DP-GEN, DeepH, DeePTB, PyATB, Hefei-NAMD, PEXSI, etc., further expanding its functionality and applicability in multi-scale calculations and electronic structure analysis.
@@
In conclusion, with the support of AI techniques, ABACUS has emerged as a robust and flexible platform for electronic structure calculations and analysis, well-positioned to address the challenges in materials science, chemistry, and physics within the AI era. Its ongoing development and integration with AI technologies are expected to further enhance its capabilities, strengthening its role as a valuable tool for researchers and scientists in the field.
@@ Acknowledgements.
We thank Han Wang and Weinan E for many helpful discussions. M.C. gratefully acknowledges AI for Science Institute, Beijing (AISI), as well as funding support from the National Key R&D Program of China under Grant No.2025YFB3003603 and the National Natural Science Foundation of China (grant no. 11988102, 12135002, 12122401,12074007). S.X. gratefully acknowledges funding support from the National Natural Science Foundation of China (grant no. 52273223), DP Technology Corporation (grant no. 2021110016001141), School of Materials Science and Engineering at Peking University, and the AI for Science Institute, Beijing (AISI). S.L. acknowledges the supports from National Natural Science Foundation of China (92370104) and Natural Science Foundation of Zhejiang Province (2022XHSJJ006). The computational resource is provided by Westlake HPC Center. M.L. gratefully acknowledges the financial support of the National Natural Science Foundation of China (No. 52276212), National Key Research and Development Program of China (No. 2022YFB3803600). Q.O. acknowledges research grants from China Petroleum &\& Chemical Corp (funding number 124014). T.W. acknowledges support by The University of Hong Kong (HKU) via seed funds (2201100392, 2409100597). Y.X. acknowledges support by the National Key Basic Research and Development Program of China (grant nos. 2024YFA1409100 and 2023YFA1406400), the National Natural Science Foundation of China (grants nos. 12334003, 12421004, and 12361141826), and the National Science Fund for Distinguished Young Scholars (grant no. 12025405). W.Z. gratefully acknowledges AI for Science Institute, Beijing (AISI), as well as support from the Hongyi postdoctoral fellowship of Wuhan University.